### v1

@@ Abstract
Context:We track and investigate from white-light data taken with the Wide-field Instrument for Solar PRobe (WISPR) aboard Parker Solar Probe (PSP), localized density enhancements, reflecting small-scale magnetic structures belonging to a filament-related coronal mass ejection (CME).
Aims:We aim to investigate the 3D location, morphology and evolution of the internal magnetic fine structures of CMEs. Specifically, we ask: what is the physical origin of the small-scales in the WISPR images? How do these structures evolve over time? What is their relationship with the filament/source region and the flux rope?
Methods:The fast tangential motion of the PSP spacecraft during its perihelion permits viewing the same event from multiple angles in short times relative to the event’s evolution. Hence, we can derive the three-dimensional information of selected CME features from a single spacecraft using triangulation techniques.
Results: We group small-scale structures with roughly similar speeds, longitude and latitude, into three distinct morphological groups. We find twisted magnetic field patterns close to the eastern leg of the CME that may be related to ’horns’ outlining the edges of the flux-rope cavity. Aligned thread-like bundles are identified close to the western leg. They may be related to confined density enhancements evolving during the filament eruption. High density blob-like features (magnetic islands) are widely spread in longitude ($\sim$40°) close to the flanks and rear part of the CME. We also note that the large-scale outer envelope of the CME, seen clearly from 1 AU, is not well observed in PSP.
Conclusions:We demonstrate that CME flux ropes may comprise different morphological groups with a cluster behavior, apart from the blobs which instead span a wide range of longitudes. This may hint either to the three-dimensionality of the post-CME current sheet (CS) or to the influence of the ambient corona in the evolutionary behavior of the CS. Importantly, we show that the global appearance of the CME can be very different in WISPR (0.11–0.16 AU) and instruments near 1 AU because of shorter line-of-sight integration of WISPR.
Internal magnetic field structures observed by PSP/WISPR in a filament related coronal mass ejection
G.M. Cappello 11    M. Temmer 11    A. Vourlidas 22    C. Braga 22    P.C. Liewer 33    J. Qiu 44
G. Stenborg
22
   A. Kouloumvakos 22    A.M. Veronig 1155    V. Bothmer 66
(Received February 14, 2024; accepted )
@@ 1 Introduction
Coronal mass ejections (CMEs) refer to the substantial expulsion of plasma and magnetic flux from the solar corona into interplanetary space. Their driver is a magnetic flux rope structure, which may not necessarily contain cold filamentary material (e.g., Chen, 2011), accelerated outward by magnetic reconnection through the Lorentz force (Forbes, 2000; Vršnak, 2001). Using multiple viewpoints of CMEs through stereoscopic coronagraph images from the Solar TErrestrial RElations Observatory (STEREO, Kaiser et al., 2008), we have gained valuable insights about their three-dimensionality when observed from 1 AU. Typically, but not always, we observe a CME morphology that includes a faint front (a shock sheath if the CME is fast enough) followed by a bright loop-like leading edge, and then a dim region corresponding to a low plasma-beta structure delineating the flux rope (Vourlidas et al., 2013; Temmer & Bothmer, 2022). When interacting with the near-Earth environment, CMEs can drive a variety of space weather effects, such as geomagnetic storms, disruption of satellite communication systems or ground-induced currents (see reviews by e.g., Pulkkinen, 2007; Gopalswamy, 2022). Hence, the evolution of CMEs from Sun to Earth has received particular attention over the years. Several recent reviews give an overview of our current understanding of the physics of CME evolution, their interaction with the ambient solar wind, state-of-the-art modeling and analysis methods, and the gaps that we need to fill in the future (e.g., Green et al., 2018; Temmer, 2021; Zhang et al., 2021; Gopalswamy, 2022; Mishra & Teriaca, 2023; Temmer et al., 2023).
@@
High-resolution observations indicate that filaments are composed of a collection of thin threads situated above the photospheric magnetic neutral line (e.g., Tandberg-Hanssen, 1995; Aulanier & Schmieder, 2002; Okamoto et al., 2007). Horizontal threads may be associated with cool material suspended between the dips of a sheared arcade or twisted flux rope, while vertical threads are not yet well understood (see e.g., Su et al., 2015; Guo et al., 2021). Schmieder et al. (2014) and Ruan et al. (2018) argued that vertical threads are due to the accumulation of many small dips containing short threads.
@@
Although small-scale CME structures have been studied in the past, the accessible scales depend on the resolution of the imaging instrument at hand. Every new generation of white-light instruments reveals ever finer details about the internal structures of CMEs and their evolution (Howard et al., 2023). Here, we investigate a variety of transient density enhancements to trace the magnetic fine structure of a filament-related CME observed by the Wide-Field Imager for Solar PRobe (WISPR; Vourlidas et al., 2016) onboard the Parker Solar Probe (PSP; Fox et al., 2016).
@@
WISPR provides stunning views of the off-limb corona with higher sensitivity compared to coronagraphic imagery from missions located near 1 AU, such as the Solar and Heliospheric Observatory (SOHO, see Brueckner et al., 1995), Domingo et al. (1995) and STEREO (see Kaiser et al., 2008; Howard et al., 2008). Illing & Hundhausen (1983) first identified looplike structures, visible within the ejecta, as bright coronal material in the form of a concave upward arch in images from the Solar Maximum Mission (SMM; Bohlin et al., 1980). These fine structures, namely the ’horns’, were also observed by Vourlidas et al. (2013) in SOHO coronagraphic images. They highlight the magnetic flux rope (MFR) nature of the cavity. Webb & Cliver (1995) and Webb et al. (2003) performed statistical studies on SMM data on concave-outward regions, finding that about half of them were followed by coaxial and bright rays, suggesting the formation of the current sheets. The latter, usually depicted as a 2D surface, may last several hours and extends in the outer corona. Post-CME blobs are usually related to the post-CME current sheet and plasmoid instabilities such as magnetic islands(see, e.g., Ciaravella & Raymond, 2008; Schanche et al., 2016; Webb & Vourlidas, 2016; Lee et al., 2020). Many simulations were performed to understand the nature of these blobs. For example, Riley et al. (2007) considered bursty reconnection in the current sheet, as modeled by a MHD simulation, and suggested the formation of blobs and their acceleration along the rays through the energy released during the reconnection process. Also, Poirier et al. (2023) modeled blobs by simulating an intermittent release of density structures generated by tearing-induced magnetic reconnection at the tip of helmet streamers and propagating in the slow solar wind.
@@
WISPR has already made close-up views of CMEs, e.g., for distances of $\sim$0.062 AU during Encounter 10 (E10). (Howard et al., 2022) highlighted the presence of brightness inhomogeneities that appear as blobs and smaller MFR structures within the larger MFR. Patel et al. (2023) studied an arc-shaped concave-up structure in a WISPR CME. In addition, PSP traversed CMEs capturing both remote sensing data and in-situ measurements of the plasma properties and the magnetic field. (Romeo et al., 2023) reported magnetic field variations within a CME containing enhancements, small rotations, and turbulence. Other highlights of the PSP mission are summarized in Raouafi et al. (2023).
@@
Usually 3D reconstructions on the global shape of the CME are performed based on multi-viewpoint images from spacecraft located at 1 AU. From these, the CMEs’ propagation direction, angular width, and tilt are derived (Thernisien et al., 2009; Kouloumvakos et al., 2022). Another method is to obtain the CME direction and speed from elongation measurements of well-observed CME fronts using single spacecraft methods such as Point-P and fixed-$\phi$ methods applied on heliospheric images (see e.g., Kahler & Webb, 2007; Wood et al., 2009; Lugaz et al., 2009). However, a recent case study by Patel et al. (2023) yields unrealistic speed results, demonstrating the shortcomings of traditional approaches for deriving kinematics when imaging from close proximity to the transient.
@@
A better approach to derive reliable kinematics and directions of coronal structures or CMEs is the geometric triangulation method developed specifically for WISPR by Liewer et al. (2019, 2020). It leverages the fast motion of the PSP spacecraft and uses the observations of the same event from multiple angles in a short time to perform triangulation with a single spacecraft. This allows us to track and investigate individual small-scale structures, which are part of the global CME. The technique has already been successfully applied in PSP studies to obtain the characteristics of CMEs and other coronal structures (Liewer et al., 2020, 2021, 2022), as well as the deflection (Braga & Vourlidas, 2021) and deformation of the CME frontal structures (Braga et al., 2022).
@@
In this study, we investigate the evolution and morphology of small-scale magnetic field structures observed inside a CME, and try to understand their relation to the associated erupted filament imaged by EUV instruments from 1 AU. How do these structures form and develop? How are they related to the larger-scale MFR in which they are embedded? The proximity of WISPR further allows us to study the effect of projection and line-of-sight integration on the CME intensity and to investigate where the small-scale structures are located with respect to the global appearance of the CME.
@@
This paper is organized as follows. In Section 2, we describe the event and observational data used in this study. Section 3 presents the methods used to investigate the small-scale structures observed during the passage of the CME with observations from PSP and also covers the CME 3D global reconstruction. In Section 4 we present our results for the morphology and 3D information, which are discussed and summarized in Section 5. The Appendix A covers a more detailed description of the geometrical triangulation method applied to PSP/WISPR observations.
@@ 2 Event description and data sources
Figure 1: Overview of the December 8, 2022 event showing the spacecraft constellation, using a Heliocentric inertial (HCI) coordinate system, together with white-light imaging data (with EUV inserts from SUVI in 304 Å, and EUVI in 195 Å) recorded from the various spacecraft vantage points (STA/COR2 in panel (B), SOHO/LASCO/C3 in panel (C), and PSP/WISPR in panel (D)). The red arrows in panels (C) and (D) highlight the localized density enhancements, referring to the magnetic fine structures contained in the core of the CME (see Section 4.4). The spacecraft constellation shows PSP orbit for E14 delineated with the blue dashed line, and the PSP trajectory during the event with the blue solid line. PSP longitude changes of $\sim 24^{\circ}$ over the entire duration of the event (15:30 UT on December 8 to 23:30 UT on December 9). The FoV of WISPR-I (red solid line) and WISPR-O (red dashed line) are also shown together with the CME propagation direction and width, obtained using the GCS reconstruction results (see Section 4.3), at four different instants of time (16:30 UT [0.11 AU], 19:30 UT [0.17 AU] and 23:22 UT [0.23 AU] on December 8 and 16:30 UT [0.51 AU] on December 9).
@@
We investigate the CME event that occurred on December 8, 2022, associated with an eruptive filament over the southwestern limb of the Sun (from Earth view). The first CME front appeared in the LASCO/C2 field of view at 04:12 UT and in C3 at 10:42 UT (at a position angle of 240°). According to the SOHO/LASCO CME catalog [^1] https://cdaw.gsfc.nasa.gov/CME_list/ the CME had a projected linear speed of $\approx$ 300 km/s and an angular width of 100°.
@@
For the event analysis, we used data from various viewpoints and instruments covering the white-light and EUV wavelength ranges. White-light data are obtained from the Solar and Heliospheric Observatory (SoHO; Domingo et al., 1995), located at the Lagrangian point L1, operating two white-light coronagraphs, LASCO/C2 with a field of view (FoV) of 1.5–6 Rdirect-$\odot$ and LASCO/C3 with a FoV of 3.7–32 Rdirect-$\odot$. The Sun-Earth Connection Coronal and Heliospheric Investigation (SECCHI; Howard et al., 2008) aboard the STEREO-A (STA) spacecraft provides a different perspective. STA observed the event from about 1 AU and 14° West of the Sun-Earth line. We used data from the SECCHI instrument suite including the inner white-light coronagraph COR1 (FoV: 1.5–4 Rdirect-$\odot$) and the outer COR2 (FoV: 2.5–15 Rdirect-$\odot$) as well as white-light heliospheric imagers, HI-1 and HI-2, covering a FoV up to Earth location and beyond.
@@
Our main focus are the WISPR data, providing unprecedented remote sensing white-light images from a close-in orbit. WISPR comprises two telescopes. The inner one, WISPR-I images an elongation of ${}^{\circ}$${}^{\circ}$, while the outer, WISPR-O, covers ${}^{\circ}$${}^{\circ}$. For more details on the WISPR instrument, we refer to Vourlidas et al. (2016).
@@
Due to its highly elliptical orbit, the PSP heliocentric distance and view angle change drastically during its orbit around the Sun. Therefore, the WISPR FoVs and temporal cadences change accordingly. The event under study was observed during Encounter 14 (E14). We use observations from December 8, 2022 15 UT to December 9, 2022 23:30 UT. During that time PSP moved from ${}^{\circ}$ (0.16 AU) to ${}^{\circ}$ (0.11 AU) in longitude (heliocentric distance) in the Heliocentric Inertial (HCI) coordinate system.
@@
The time cadence increased from $\sim$30 min and $\sim$5 min as the spacecraft approached the Sun during the event. Furthermore, the event was captured in the ’wave turbulence’ mode (WT), where a subfield in WISPR-I is captured at 1 min cadence for several hours. WT data are available from 08:03 UT to 20:53 UT on December 8, 2022. No WISPR-O images are acquired during the WT mode, but WISPR-I full FoV images are available. WT image data allow us to follow the small-scale structures with less ambiguity and to better appreciate the complexity of the CME. We make use of the ’LW’-processed WISPR images for the analysis (see the Appendix in Howard et al., 2022, for more details). The ’LW’ technique increases the contrast and facilitates the tracking of the features.
@@
To relate the features observed in white-light to the solar source region, we use available EUV data in the wavelength ranges 171 Å, 195 Å and 304 Å from the Extreme Ultraviolet Imager (EUVI) onboard STA with a FoV up to ${}_{\odot}$ (Wuelser et al., 2004). The location of the erupting prominence, being close to the solar limb, allows us to track the event back to the solar surface. For the Earth perspective, we used EUV data from the Solar Ultraviolet Imager (SUVI; Darnel et al., 2022) onboard the Geostationary Operational Environmental Satellite(GOES-R; Krimchansky et al., 2004), which has a FoV similar to EUVI. We note that we do not discuss the initiation and triggering of the filament eruption itself.
@@
The left panel of Fig. 1 provides an overview of the spacecraft configuration on December 8, 2022 and marks the CME trajectory at different times. The right panels in Fig. 1 show combined COR2/EUVI-195Å  (panel B), C3/SUVI-304Å  (panel C), and WISPR-I/O composite snapshots (panel D). The COR2 snapshot was obtained at 16:23 UT. The C3 images were taken at 16:23 UT, 19:23 UT, and 23:38 UT and reveal aligned fine structures marked by a red arrow. The WISPR composite, taken at around 22:15 UT (or $\sim 22.22$ UT in Earth time), shows an incredible amount of detail within the CME, marked with a red arrow as well.

### v2

@@ Abstract
Context. We investigated the coronal mass ejection (CME) related to an eruptive filament over the southwestern solar limb on December 8, 2022, at around 8 UT. We tracked localized density enhancements reflecting the magnetic structures using white-light data taken with the Wide-field Instrument for Solar PRobe (WISPR) aboard the Parker Solar Probe (PSP).
Aims. We aim to investigate the 3D location, morphology and evolution of the internal magnetic fine structures of CMEs. Specifically, we focused on the physical origin of the features in the WISPR images, how the white-light structures evolve over time, and their relationship with the source region, filament, and the flux rope.
Methods. The fast tangential motion of the PSP spacecraft during its perihelion permits a single event to be viewed from multiple angles in short times relative to the event’s evolution. Hence, three-dimensional information of selected CME features can be derived from this single spacecraft using triangulation techniques.
Results. We grouped small-scale structures with roughly similar speeds, longitude, and latitude into three distinct morphological groups. We found twisted magnetic field patterns close to the eastern leg of the CME that may be related to ”horns” outlining the edges of the flux-rope cavity. We identified aligned thread-like bundles close to the western leg, and they may be related to confined density enhancements evolving during the filament eruption. High density blob-like features (magnetic islands) are widely spread in longitude ($\sim$40°) close to the flanks and the rear part of the CME. We also note that the large-scale outer envelope of the CME, seen clearly from 1 AU, was not well observed by PSP.
Conclusions. We demonstrate that CME flux ropes, apart from the blobs, may comprise different morphological groups with a cluster behavior; the blobs instead span a wide range of longitudes. This finding may hint at either the three-dimensionality of the post-CME current sheet (CS) or the influence of the ambient corona in the evolutionary behavior of the CS. Importantly, we show that the global appearance of the CME can be very different in WISPR (0.11–0.16 AU) and the instruments near 1 AU because of the shorter line-of-sight integration of WISPR.
@@ 1 Introduction
Coronal mass ejections (CMEs) refer to the substantial expulsion of plasma and magnetic flux from the solar corona into interplanetary space. Their driver is a magnetic flux rope structure, which may not necessarily contain cold filamentary material (e.g., Chen, 2011), accelerated outward by magnetic reconnection through the Lorentz force (Forbes, 2000; Vršnak, 2001). Using multiple viewpoints of CMEs through stereoscopic coronagraph images from the Solar TErrestrial RElations Observatory (STEREO; Kaiser et al., 2008), valuable insights have been gained about their three-dimensionality when observed from 1 AU. Typically, but not always, a CME morphology is observed that includes a faint front (a shock sheath if the CME is fast enough) followed by a bright loop-like leading edge and then a dim region corresponding to a low plasma-beta structure delineating the flux rope (Vourlidas et al., 2013; Temmer & Bothmer, 2022). When interacting with the near-Earth environment, CMEs can drive a variety of space weather effects, such as geomagnetic storms, disruption of satellite communication systems, and ground-induced currents (see reviews by e.g., Pulkkinen, 2007; Gopalswamy, 2022). Hence, the evolution of CMEs from the Sun to Earth has received particular attention over the past years. Several recent reviews have given an overview of our current understanding of the physics of CME evolution, their interaction with the ambient solar wind, the state of the art in modeling and analysis methods, and the gaps that need to be filled in the future (e.g., Green et al., 2018; Temmer, 2021; Zhang et al., 2021; Gopalswamy, 2022; Mishra & Teriaca, 2023; Temmer et al., 2023).
@@
High-resolution observations have indicated that filaments are composed of a collection of thin threads situated above the photospheric magnetic neutral line (e.g., Tandberg-Hanssen, 1995; Aulanier & Schmieder, 2002; Okamoto et al., 2007). Horizontal threads may be associated with cool material suspended between the dips of a sheared arcade or twisted flux rope, while vertical threads are not yet well understood (see e.g., Su et al., 2015; Guo et al., 2021). Schmieder et al. (2014) and Ruan et al. (2018) have argued that vertical threads are caused by the accumulation of many small dips containing short threads.
@@
Although small-scale CME structures have been studied in the past, the accessible scales depend on the resolution of the imaging instrument at hand. Every new generation of white-light instruments reveals ever-finer details about the internal structures of CMEs and their evolution (Howard et al., 2023). Here, we investigate a variety of transient density enhancements in order to trace the magnetic fine structure of a filament-related CME observed by the Wide-Field Imager for Solar PRobe (WISPR; Vourlidas et al., 2016) on board the Parker Solar Probe (PSP; Fox et al., 2016).
@@
WISPR provides stunning views of the off-limb corona with a higher sensitivity compared to coronagraphic imagery from missions located near 1 AU, such as the Solar and Heliospheric Observatory (SOHO; see Brueckner et al., 1995; Domingo et al., 1995) and STEREO (see Kaiser et al., 2008; Howard et al., 2008). Illing & Hundhausen (1983) first identified loop-like structures visible within the ejecta as bright coronal material in the form of a concave upward arch in images from the Solar Maximum Mission (SMM; Bohlin et al., 1980). These fine structures, namely the ’horns’, were also observed by Vourlidas et al. (2013) in SOHO coronagraphic images. The structures highlight the magnetic flux rope (MFR) nature of the cavity. Webb & Cliver (1995) and Webb et al. (2003) performed statistical studies on SMM data of concave-outward regions, and they found that about half of them were followed by coaxial and bright rays, suggesting the formation of the current sheets. The latter, usually depicted as a 2D surface, may last several hours and extends to the outer corona. Post-CME blobs are usually related to the post-CME current sheet and plasmoid instabilities such as magnetic islands(see, e.g., Ciaravella & Raymond, 2008; Schanche et al., 2016; Webb & Vourlidas, 2016; Lee et al., 2020). Many simulations have been performed to understand the nature of these blobs. For example, Riley et al. (2007) considered bursty reconnection in the current sheet, as modeled by a MHD simulation, and suggested the formation of blobs and their acceleration along the rays to be through the energy released during the reconnection process. Also, Poirier et al. (2023) modeled blobs by simulating an intermittent release of density structures generated by tearing-induced magnetic reconnection at the tip of helmet streamers and propagating in the slow solar wind.
@@
WISPR has already made close-up views of CMEs (e.g., for distances of $\sim$0.062 AU during Encounter 10, or E10). (Howard et al., 2022) highlighted the presence of brightness inhomogeneities that appear as blobs and smaller MFR structures within the larger MFR. Patel et al. (2023) studied an arc-shaped concave-up structure in a WISPR CME. In addition, PSP has traversed CMEs capturing both remote sensing data and in situ measurements of the plasma properties and the magnetic field. (Romeo et al., 2023) reported magnetic field variations within a CME containing enhancements, small rotations, and turbulence. Other highlights of the PSP mission are summarized in Raouafi et al. (2023).
@@
Usually, 3D reconstructions of the global shape of a CME are performed based on multi-viewpoint images from spacecraft located at 1 AU. From these, a CME’s propagation direction, angular width, and tilt are derived (Thernisien et al., 2009; Kouloumvakos et al., 2022). Another method is to obtain the CME direction and speed from elongation measurements of well-observed CME fronts using single spacecraft methods, such as Point-P and fixed-$\phi$ methods, applied to heliospheric images (see e.g., Kahler & Webb, 2007; Wood et al., 2009; Lugaz et al., 2009). However, a recent case study by Patel et al. (2023) yielded unrealistic speed results, demonstrating the shortcomings of traditional approaches for deriving kinematics when imaging in proximity to a transient.
@@
A better approach to derive reliable kinematics and directions of coronal structures or CMEs is the geometric triangulation method developed specifically for WISPR by Liewer et al. (2019, 2020). It leverages the fast motion of the PSP spacecraft and uses observations of a single event from multiple angles across a short time frame to perform triangulation with only a single spacecraft. This allowed us to track and investigate individual small-scale structures, that are part of the global CME. The technique has already been successfully applied in PSP studies to obtain the characteristics of CMEs and other coronal structures (Liewer et al., 2020, 2021, 2022), as well as the deflection (Braga & Vourlidas, 2021) and deformation of CME frontal structures (Braga et al., 2022).
@@
In this study, we investigate the evolution and morphology of small-scale magnetic field structures observed inside a CME, and we try to understand their relation to the associated erupted filament imaged by EUV instruments from 1 AU. Specifically, we are concerned with how these structures form and develop and how they related to the larger-scale MFR in which they are embedded. The proximity of WISPR to the Sun further allowed us to study the effect of projection and line-of-sight integration on the CME intensity and to investigate where the small-scale structures are located with respect to the global appearance of the CME.
@@
This paper is organized as follows. In Sect.  2, we describe the event and observational data used in this study. Section  3 presents the methods used to investigate the small-scale structures observed during the passage of the CME with observations from PSP and also covers the CME 3D global reconstruction. In Sect.  4 we present our results for the morphology and 3D information, which are discussed and summarized in Sect.  5. Finally, Appendix A covers a more detailed description of the geometrical triangulation method applied to the PSP/WISPR observations.
@@ 2 Event description and data sources
Figure 1: Overview of the December 8, 2022 event showing the spacecraft constellation (panel A), using an HCI coordinate system, together with white-light imaging data (with EUV inserts from SUVI in 304 Å, and EUVI in 195 Å) recorded from the various spacecraft vantage points. Panel B shows the vantage point from STA/COR2, panel C shows the vantage point from SOHO/LASCO/C3 and panel D, shows the vantage point from PSP/WISPR. The red arrows in panels (C) and (D) highlight the localized density enhancements, referring to the magnetic fine structures contained in the core of the CME (see Sect.  4.4). The spacecraft constellation shows PSP’s orbit for E14 with the blue dashed line and the PSP trajectory during the event with the blue solid line. The PSP longitude changed $\sim 24^{\circ}$ over the entire duration of the event (15:30 UT on December 8 to 23:30 UT on December 9). The FoV of WISPR-I (red solid line) and WISPR-O (red dashed line) are also shown together with the CME propagation direction and width, which was obtained using the GCS reconstruction results (see Sect.  4.3), at four different instants of time (16:30 UT [0.11 AU], 19:30 UT [0.17 AU], and 23:22 UT [0.23 AU] on December 8 and 16:30 UT [0.51 AU] on December 9).
@@
We investigated the CME event that occurred on December 8, 2022, and is associated with an eruptive filament over the southwestern limb of the Sun (from Earth view). The first CME front appeared in the LASCO/C2 field of view (FoV) at 04:12 UT and in C3 at 10:42 UT (at a position angle of 240°). According to the SOHO/LASCO CME catalog [^1] https://cdaw.gsfc.nasa.gov/CME_list/, the CME had a projected linear speed of $\approx$ 300 km/s and an angular width of 100°.
@@
For the event analysis, we used data from various viewpoints and instruments covering the white-light and EUV wavelength ranges. White-light data were obtained from the SOHO observatory Domingo et al. (1995) located at the Lagrangian point L1 and operating two white-light coronagraphs: LASCO/C2, with an field of view (FoV) of 1.5–6 Rdirect-$\odot$, and LASCO/C3, with an FoV of 3.7–32 Rdirect-$\odot$. The Sun-Earth Connection Coronal and Heliospheric Investigation (SECCHI; Howard et al., 2008) aboard the STEREO-A (STA) spacecraft provides a different perspective, as STA observed the event from about 1 AU and 14° west of the Sun-Earth line. We used data from the SECCHI instrument suite, including the inner white-light coronagraph COR1 (FoV: 1.5–4 Rdirect-$\odot$) and the outer COR2 (FoV: 2.5–15 Rdirect-$\odot$), as well as white-light heliospheric imagers, HI-1 and HI-2, covering an FoV up to Earth location and beyond.
@@
Our main focus is the WISPR data, as they provide unprecedented remote sensing white-light images from a close-in orbit. WISPR comprises two telescopes. The inner one, WISPR-I, images an elongation of 13.5–53.0, while the outer, WISPR-O, covers 50.5–108.5. For more details on the WISPR instrument, we refer to Vourlidas et al. (2016).
@@
Due to the PSP highly elliptical orbit, its heliocentric distance and its viewing angle changes drastically during its orbit around the Sun. Therefore, the WISPR FoVs and temporal cadences change accordingly. The event under study was observed during Encounter 14 (E14). We used observations from December 8, 2022, 15 UT to December 9, 2022, 23:30 UT. During that time, PSP moved from --41.9 (0.16 AU) to --18.2 (0.11 AU) in longitude (heliocentric distance) in the heliocentric Inertial (HCI) coordinate system.
@@
The time cadence increased from $\sim$30 min and $\sim$5 min as the spacecraft approached the Sun during the event. Furthermore, the event was captured in the ”wave turbulence” mode (WT), where a subfield in WISPR-I is captured at 1-min cadence for several hours. The WT data from 08:03 UT to 20:53 UT on December 8, 2022, are available. No WISPR-O images were acquired during the WT mode, but WISPR-I full FoV images were available. The WT image data allowed us to follow the small-scale structures with less ambiguity and to better appreciate the complexity of the CME. We made use of the ”LW”-processed WISPR images for the analysis (see the appendix in Howard et al. (2022), for more details). The LW technique increases the contrast and facilitates the tracking of the features.
@@
To relate the features observed in white-light to the solar source region, we used the available EUV data in the wavelength ranges of 171 Å , 195 Å , and 304 Å  from the Extreme Ultraviolet Imager (EUVI) on board STA with an FoV up to 1.7R (Wuelser et al., 2004). As it was close to the solar limb, the location of the erupting promience allowed us to track the event back to the solar surface. For the Earth perspective, we used EUV data from the Solar Ultraviolet Imager (SUVI; Darnel et al., 2022) onboard the Geostationary Operational Environmental Satellite(GOES-R; Krimchansky et al., 2004), which has an FoV similar to EUVI. We note that we do not discuss the initiation and triggering of the filament eruption itself.
@@
The left panel of Fig.  1 provides an overview of the spacecraft configuration on December 8, 2022, and marks the CME trajectory at different times. The right panels in Fig.  1 show combined COR2/EUVI-195Å  (panel B), C3/SUVI-304Å  (panel C), and WISPR-I/O composite snapshots (panel D). The COR2 snapshot was obtained at 16:23 UT. The C3 images were taken at 16:23 UT, 19:23 UT, and 23:38 UT, and they reveal aligned fine structures (marked by a red arrow in panel C). The WISPR composite, taken at around 22:15 UT (or $\sim 22.22$ UT in Earth time), shows an incredible amount of detail within the CME, marked with a red arrow in panel D.